• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name M Friend

    Log entry time 20:16:01 on November 14, 2009

    Entry number 299896

    This entry is a followup to: 299892

    keyword=Compton change

    The shielding helped the base current problem, but did not fix it completely. Instead of fluctuating from 972uA to 1050uA with beam trips (100uA beam), the base current is fluctuating from 972uA to 984uA. During HAPEX the base current did not fluctuate at all. Perhaps we need even more shielding.

    A copy of this log entry has been emailed to: rom,riordan,bquinn@cmu.edu,kdp2c@virginia.edu